# Capacity Gains in MIMO Systems with Few-Bit ADCs Using Nonlinear Analog Circuits

### v2

@@ VI A Hybrid Beamforming Architecture with One-bit ADCs
In the previous sections, we have investigated the channel capacity for SISO systems (under analog beamforming) equipped with different collections of implementable analog functions. In this section, we consider hybrid beamforming with one-bit ADCs, where the beamforming vector at the receiver $\mathbf{w}\in\mathbb{R}^{n_{r}\times n_{s}}$ maps the received signal $Y^{n_{r}}$ to $\widetilde{Y}^{n_{s}}$, where $n_{s}>1$ (Figure 2). In this case, we provide a quantization setup, using envelope detectors, which accommodates QAM modulation, and derive an inner bounds to the system capacity. In the next section, we numerically evaluate the resulting capacity and provide comparisons with prior works.
@@ VI-A Quantizer Construction
We assume that $n_{q}>n_{s}$, otherwise, one can use analog beamformers to further reduce the dimension of the beamformer output without performance loss in terms of achievable rates. As mentioned in Section II-A a quantizer is characterized by its analog processing functions $f_{j}(\cdot),j\in[n_{q}]$ and one-bit ADC thresholds $t^{n_{q}}\in\mathbb{R}^{n_{q}}$. Let us fix a threshold step parameter $\zeta>0$. We take the analog processing functions as follows:
$f_{j}(\widetilde{y}^{n_{s}})=\begin{cases}\widetilde{y}_{j}\quad&\text{ if }j\leq n_{s},\\ |\widetilde{y}_{\bar{j}}|&\text{ if }j>n_{s},\end{cases}\qquad$
where $\bar{j}$ is the module $n_{s}$ residual of $j$. We take the threshold values as follows:
$t_{j}=\begin{cases}0\quad&\text{ if }j\leq n_{s},\\ \lfloor\frac{j}{n_{q}}\rfloor\zeta&\text{ if }j>n_{s}.\end{cases}\qquad$
@@
Figure 7: The quantizer outputs and Voronoi regions in Example 5.
@@
Let $n_{s}=2$, $n_{q}=6$, and $\zeta=1$. Then, for the construction described above, the six one-bit ADC operations are as follows:
$Q_{1}:\widetilde{y}_{1}\lessgtr 0,\quad Q_{2}:\widetilde{y}_{2}\lessgtr 0,\quad Q_{3}:|\widetilde{y}_{1}|\lessgtr 1,\quad$
$Q_{4}:|\widetilde{y}_{2}|\lessgtr 1,\quad Q_{5}:|\widetilde{y}_{1}|\lessgtr 2,\quad Q_{6}:|\widetilde{y}_{2}|\lessgtr 2.$
@@ VI-B Achievable Rates at High SNR
As argued in Section IV, the high SNR capacity is equal to the maximum number of quantization regions which can be generated given the number of ADCs $n_{q}$, with ADC resolution equal to $\ell$, and set of implementable analog functions $\mathcal{F}_{a}$. The following theorem provides upper and lower bounds on the high SNR channel capacity of beamforming architectures equipped with envelope detectors for analog signal processing.
Theorem 5 (High SNR Capacity of Hybrid-Beamforming System).
Assume that the channel matrix observed after beamforming is full-rank, i.e. the rank of $\mathbf{w}^{H}\mathbf{h}\mathbf{f}$ is equal to $n_{s}$. Let $C_{env}(P_{T},P_{ADC},n_{s},\delta,\ell)$ denote the channel capacity under power constrain $P_{T}$. Then,
$n_{s}\left(1+\log{(\ell-1)}+\log\left(\frac{n_{q}-n_{s}}{n_{s}}\right)\right)\leq\lim_{P_{T}\to\infty}$ (19)
$\qquad\qquad\qquad C_{env}(P_{T},P_{ADC},n_{s},1)\leq\log\sum_{k=0}^{n_{s}}{2(\ell-1)n_{q}\choose k},$
where $n_{q}=\frac{P_{ADC}}{2\alpha}$.
The lower bound in Equation (19) is achieved by the quantizer described in this section. To see this, note that by construction, the quantizer partitions each axis into $2(\ell-1)\left(\frac{n_{q}-n_{s}}{n_{s}}\right)$ intervals, and each resulting quantization region is mapped to a unique quantizer output (e.g., Figure 7). So, the total number of unique quantizer outputs is $|\mathcal{C}|=(2(\ell-1)(\frac{n_{q}-n_{s}}{n_{s}}))^{n_{s}}$. The result follows by noting that the communication rate is $\log{|\mathcal{C}|}$. The upper bound follows by counting the number of partition regions generated by $2(\ell-1)n_{q}$ hyperplanes in general position in the $n_{s}$-dimensional Euclidean space (e.g. [30]). Figure 8 provides numerical simulations of the i) upper bounds and ii) lower bounds in Equation (19) and iii) the high SNR channel capacity under hybrid beamforming without analog processing derived in [30] for $n_{s}=3$. It can be observed that the proposed architecture outperforms the one in [30] if the number of one-bit ADCs is larger than $n_{q}=8$.
Figure 8: Channel capacity for i) the proposed architecture and $n_{q}$ one-bit ADCs (lower-bound in Theorem 5), ii) no analog processing prior to quantization and $2n_{q}$ one-bit ADCs (upper-bound in Theorem 5) and iii) no analog processing prior to quantization and $n_{q}$ one-bit ADCs ([30]).
@@ VII Numerical Analysis of Channel Capacity
In this section, we provide a numerical analysis of the capacity bounds derived in the prequel and evaluate the gains due to the use of nonlinear analog components in the receiver terminal.
VII-A Capacity Evaluation for Envelope Detector Architectures
@@
We compute an inner-bound to the capacity expression of the SISO system in Theorem 3, for various SNR values and as a function of the number of ADCs $n_{q}$ and the number of concatenated envelope detectors $\delta$. To this end, we first use the extension of the Blahut-Arimoto algorithm to discrete memoryless channels with input cost constraints given in [52] to find the best input distribution. Then, we conduct a brute-force search over all possible uniform quantizers. In order to find the mass points of $X$, we discretize the real-line using a grid with step-size 0.1, and optimize the distribution over the resulting discrete space. Fig. 9 shows the resulting achievable rates for SNRs in the range of -5 to 40 dB for various values of $(n_{q},\delta)$. Observe that without nonlinear analog processing, for $\ell=2$, the high SNR capacity is $\log{n_{q}+1}$ [30]. So, for instance, for $n_{q}=\delta=2$, the inner bound in Figure 9 surpasses the high SNR capacity without nonlinear analog processing for SNRs higher than 15dB, and its high SNR capacity is more than $25\%$ greater than the case when there is no nonlinear analog processing.
Figure 9: The set of achievable rates for various values of $(n_{q},\ell,\delta)$ for architectures using envelope detectors.
VII-B Capacity Evaluation for Polynomial Operator Architectures
We numerically evaluate the inner bound to the capacity region given in Theorem 4. Similar to the previous case, we first use the extension of the Blahut-Arimoto algorithm to find the best input distribution. Then, we conduct a brute-force search over all possible threshold vectors. To find the mass points of $X$, we discretize the real-line using a grid with step-size 0.1, and optimize the distribution over the resulting discrete space. Figure 10 shows the resulting achievable rates for SNRs in the range of -5 to 40 dB for various values of $(n_{q},\delta)$. It can be observed that the performance improvements due to the use of higher degree polynomials are more significant at high SNRs. Furthermore, it can be observed that the set of achievable rates only depends on $min(\ell^{n_{q}},(\ell-1)\delta n_{q}+1)$. As a result, for instance the achievable rate when $n_{q}=2,\ell=2,\delta=2$ is the same as that of $n_{q}=3,\ell=2,\delta=1$ as shown in the figure. So, in this case, using higher degree polynomials can compensate for a lower number of ADCs. On the other hand, the achievable rate for $n_{q}=3,\ell=2,\delta=1$ is lower than that of $n_{q}=3,\ell=2,\delta=2$ as shown in the figure. So, using higher degree polynomials leads to rate improvements in this scenario.
Figure 10: The set of achievable rates for various values of $(n_{q},\ell,\delta)$ for architectures using polynomial operators.
Figure 11: Inner-bound to achievable rates for the hybrid beamforming architecture of Section VI.
VII-C Achievable Inner-Bound for Hybrid Beamforming with Envelope Detectors
We numerically evaluate the inner-bound to the channel capacity which is achievable using the beamforming architecture described in Section VI. We have simulated the communication system for $n_{s}=2$, $\ell=2$, and $n_{q}\in\{5,6\}$ and found an estimate of the achievable rates empirically. The results are shown in Figure 11. To perform the simulation, we have optimized the threshold parameter $\zeta$, the input alphabet values $\mathbf{x}\in\mathbb{R}^{2}$, and the probability distribution $P_{X}$ on the alphabet of the input points using a gradient descent optimization method. We have simulated the channel by generating 15000 independent and identically distributed samples of noise vectors and input messages. We have used the empirical observations to estimate transition probability of the discrete channel resulting from the quantization process. We have used the Blahut-Arimoto algorithm to find the capacity of the resulting channel. It can be observed in Figure 11 that for $n_{q}=5$, the high SNR rate is larger than the $n_{s}\left(1+\log{\left(\frac{n_{q}-n_{s}}{n_{s}}\right)}\right)$ lower-bound given in Theorem 4, whereas it is equal to this lower-bound for $n_{q}=6$.
@@ VII-C Achievable Inner-Bound for Hybrid Beamforming with Envelope Detectors
Figure 12: The block diagram of a MIMO receiver using (left) conventional architectures and (right) proposed architecture.
Power Breakdown in Conventional Receiver BF LNA Phase Shifter [53] ${}^{\dagger}$ Mixer* ${}^{\ddagger}$ Total (mW) Rate Rate/Power Analog $16\times 10$ $16\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 2\times 70fJ\times 2^{8}\times 10^{9}\times 10^{3}$ 413.54 8 0.0193 Hybrid $(K=2)$ $16\times 10$ $16\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 2\times 70fJ\times 2^{8}\times 10^{9}\times 10^{3}$ 483.08 16 0.0331
Power Breakdown in Proposed Receiver (envelope detectors) BF LNA Phase Shifter [53] ${}^{\dagger}$ Mixer* NL Operator ${}^{\ddagger}$ Total (mW) Rate Rate/Power Analog $16\times 10$ $16\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 3$ $32\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ 416.54 8.807 0.0211 Hybrid $(K=2)$ $16\times 10$ $16\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 3$ $64\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ 489.08 17.615 0.0360
Power Breakdown in Proposed Receiver (2nd order polynomial) BF LNA Phase Shifter [53] ${}^{\dagger}$ Mixer* NL Operator ${}^{\ddagger}$ Total (mW) Rate Rate/Power Analog $16\times 10$ $16\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 5$ $32\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ 418.54 8.807 0.0210 Hybrid $(K=2)$ $16\times 10$ $16\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 5$ $64\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ 493.08 17.615 0.0357
Power Breakdown in Proposed Receiver (4th order polynomial) BF LNA Phase Shifter [53] ${}^{\dagger}$ Mixer* NL Operator ${}^{\ddagger}$ Total (mW) Rate Rate/Power Analog $16\times 10$ $16\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 10$ $32\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ 423.54 9.807 0.0232 Hybrid $(K=2)$ $16\times 10$ $16\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 10$ $64\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ 503.08 19.615 0.0390
@@
${}^{\dagger}$ For each RX, a four bit binary weighted current adder is implemented with 1.6mW power consumption
${}^{\ddagger}$ Both power consumption of one-bit-ADC and 8-bit-ADC are normalized from the reference
${}^{\ddagger}$ For the 4th order polynomial, the data rate and subsequently the count of ADCs is twice that of second order polynomial.
@@
Figure 13: The breakdown of power consumption for the receiver in a conventional architecture and the proposed architecture. It is evident that the proposed architecture lowers the total power consumption of the entire receiver chain substantially

### v3

@@ VI A Hybrid Beamforming Architecture with One-bit ADCs
In the previous sections, we have investigated the channel capacity for SISO systems equipped with different collections of implementable analog functions. In this section, we consider MIMO systems with one-bit ADCs. We provide a quantization setup, using envelope detectors, which accommodates QAM modulation, and derive an inner bounds to the system capacity. In the next section, we numerically evaluate the resulting capacity and provide comparisons with prior works.
@@ VI-A Quantizer Construction
We assume that $n_{q}>n_{r}$. As mentioned in Section 1 a quantizer is characterized by its analog processing functions $f_{j}(\cdot),j\in[n_{q}]$ and ADC thresholds [^5] We consider one-bit ADCs in this section, hence, each ADC has only one threshold value. $t^{n_{q}}\in\mathbb{R}^{n_{q}}$. Let us fix a threshold step parameter $\zeta>0$. We take the analog processing functions as follows:
$f_{j}({y}^{n_{r}})=\begin{cases}{y}_{j}\quad&\text{ if }j\leq n_{r},\\ |{y}_{\bar{j}}|&\text{ if }j>n_{r},\end{cases}\qquad$
where $\bar{j}$ is the modulo $n_{r}$ residual of $j$. We take the threshold values as follows:
$t_{j}=\begin{cases}0\quad&\text{ if }j\leq n_{r},\\ \lfloor\frac{j}{n_{q}}\rfloor\zeta&\text{ if }j>n_{r}.\end{cases}\qquad$ (15)
@@
Figure 5: Quantizer outputs and Voronoi regions in Example 5.
@@
Let $n_{r}=2$, $n_{q}=6$, and $\zeta=1$. Then, for the construction described above, the six one-bit ADC operations are as follows:
$q_{1}:{y}_{1}\lessgtr 0,\quad q_{2}:{y}_{2}\lessgtr 0,\quad q_{3}:|{y}_{1}|\lessgtr 1,\quad$
$q_{4}:|{y}_{2}|\lessgtr 1,\quad q_{5}:|{y}_{1}|\lessgtr 2,\quad q_{6}:|{y}_{2}|\lessgtr 2.$
@@ VI-B Achievable Rates at High SNR
As argued in Section IV, the high SNR capacity is equal to the maximum number of Voronoi regions which can be generated given the number of ADCs $n_{q}$ with $\ell$ output levels, and the set of implementable analog functions. The following theorem provides upper and lower bounds on the high SNR channel capacity of a MIMO system equipped with $n_{q}$ one-bit ADCs. The lower-bound is based on MIMO systems equipped with envelope detectors, whereas the upper-bound holds generally regardless of analog processing restrictions.
Figure 6: (a) Channel capacity for i) the proposed architecture and $n_{q}$ one-bit ADCs (lower-bound in Theorem 5), ii) no analog processing and $2n_{q}$ one-bit ADCs (upper-bound in Theorem 5) and iii) no analog processing and $n_{q}$ one-bit ADCs ( [27]). (b) Achievable rates for various $(n_{q},\delta)$ with $\ell=2$ using envelope detectors; $\delta=0$ represents the baseline without analog processing. (c) Achievable rates for the hybrid beamforming architecture of Section VI.
Theorem 5.
Given $P_{T},P_{ADC}>0$ and $n_{r},n_{q}\in\mathbb{N}$, let $C_{env}(P_{T},P_{ADC},n_{r},n_{q})$ denote the channel capacity of the MIMO system. Then,
$n_{r}\left(1+\log{(\ell-1)}+\log\left(\frac{n_{q}-n_{r}}{n_{r}}\right)\right)\!\!\leq\!\!\lim_{P_{T}\to\infty}C_{env}(P_{T},P_{ADC},n_{r},n_{q})$ (16)
$\qquad\qquad\qquad\leq\log\sum_{k=0}^{n_{r}}{2(\ell-1)n_{q}\choose k},$
where $n_{q}=\lfloor\frac{P_{ADC}}{2\alpha}\rfloor$.
The lower bound in Equation (16) is achieved by the quantizer introduced in the previous section. To see this, note that by construction, the quantizer partitions each axis into $2(\ell-1)\left(\frac{n_{q}-n_{r}}{n_{r}}\right)$ intervals, and each resulting quantization region is mapped to a unique quantizer output (e.g., Figure 5). So, the total number of unique quantizer outputs is $|\mathcal{C}|=(2(\ell-1)(\frac{n_{q}-n_{r}}{n_{r}}))^{n_{r}}$. The result follows by noting that the communication rate is $\log{|\mathcal{C}|}$. The upper bound follows by counting the number of partition regions generated by $2(\ell-1)n_{q}$ hyperplanes in general position in the $n_{r}$-dimensional Euclidean space (e.g. [27]). Figure 6(a) provides numerical simulations of the i) upper bounds and ii) lower bounds in Equation (16) and iii) the high SNR channel capacity under hybrid beamforming without analog processing derived in [27], where we have taken $n_{r}=3$. It can be observed that the proposed architecture outperforms the one in [27] if the number of one-bit ADCs is larger than $n_{q}=8$.
@@ VII Numerical Analysis of Channel Capacity
We compute an inner-bound to the capacity expression of the SISO system in Theorem 3, for various SNR values and as a function of the number of ADCs $n_{q}$ and the number of concatenated envelope detectors $\delta$. To this end, we conduct a brute-force search as described in the following. Recall that the optimization is over three sets of parameters, namely, the input points, $\mathbf{x}$, threshold vector $\mathbf{t}$, and probability distribution $P_{X}$. To search for the value of x and t, we discretize the interval $[-3,3]$ using a grid with step-size $0.1$. We search over all symmetric vectors t over the grid, where symmetry is with respect to the origin. We place each input point ${x}_{i}$ in the center of the corresponding quantization interval $[t_{i},t_{i+1}]$. Furthermore, we optimize the distribution $P_{X}$ over the resulting discrete space via the Blahut-Arimoto algorithm. That is, for each configuration, we find optimal input distribution, using the modified Blahut-Arimoto algorithm for discrete memoryless channels with input cost constraints given in [39]. Fig. 6(b) shows the achievable rates for SNRs in the range of -10 to 40 dB for various values of $(n_{q},\delta)$. Observer that for $n_{q}=\delta=2$, the inner bound in Figure 6(b) at SNRs higher than 15dB surpasses the high SNR capacity of systems without nonlinear analog processing with $n_{q}=2$ and $\delta=0$, and the high SNR capacity of the former is more than $25\%$ greater than that of the latter. Note that the achievable rate for quadratic operator architectures, for appropriately chosen parameters, would completely match the ones shown in Figure 6(b). For instance, for a polynomial operators with $\delta=2$, if the input to the ADC is $(Y-a)^{2}$ and the threshold is $b$, then the setup is equivalent to feeding $|Y-a|$ to an ADC with threshold $\sqrt{b}$. Consequently, we do not plot the resulting achievable rates using polynomial operators with $\delta=2$, since they would match the ones in Figure 6(b) for envelop detectors with $\delta=1$.
@@
Figure 7: A hybrid beamforming receiver using (left) conventional architectures and (right) proposed architecture.
Power Breakdown for Analog Beamforming with 4 Antennas
Receiver LNA Phase Shifters Combiners Mixers NL Operator ADC[40] Total (mW) Rate Rate/Power % Improvement
Linear Operator $4\times 10$ $4\times 11.5$ $1\times 19.5$ $1\times 14.2$ 00 $1\times 2\times 70fJ\times 2^{8}\times 10^{9}\times 10^{3}$ $155.54$ $8$ $0.05143$
1 Env. Detector $4\times 10$ $4\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 3$ $8\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $131.66$ $6.807$ 0.05170 0.53%
3 Env. Detectors $4\times 10$ $4\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 9$ $8\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $137.66$ $8.807$ 0.06398 24.39%
6 Env. Detectors $4\times 10$ $4\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 18$ $8\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $146.66$ $11.807$ 0.08051 56.53%
$2^{nd}$ degree Poly. $4\times 10$ $4\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 5$ $8\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $133.66$ $6.807$ 0.05093
$4^{th}$ degree Poly. $4\times 10$ $4\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 10$ $8\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $138.66$ $7.807$ 0.05631 9.47%
Power Breakdown for Analog Beamforming with 16 Antennas
Receiver LNA Phase Shifters Combiners Mixers NL Operator ADC[40] Total (mW) Rate Rate/Power % Improvement
Linear Operator $16\times 10$ $16\times 11.5$ $1\times 19.5$ $1\times 14.2$ 00 $1\times 2\times 70fJ\times 2^{8}\times 10^{9}\times 10^{3}$ $413.54$ $8$ $0.01935$
1 Env. Detector $16\times 10$ $16\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 3$ $16\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $398.62$ $7.807$ 0.01958 1.24%
3 Env. Detectors $16\times 10$ $16\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 9$ $16\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $404.62$ $9.807$ 0.02424 25.29%
6 Env. Detectors $16\times 10$ $16\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 18$ $16\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $413.62$ $12.807$ 0.03096 60.06%
$2^{nd}$ degree Poly. $16\times 10$ $16\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 5$ $16\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $400.62$ $7.807$ 0.01949 0.74%
$4^{th}$ degree Poly. $16\times 10$ $16\times 11.5$ $1\times 19.5$ $1\times 14.2$ $1\times 10$ $16\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $405.62$ $8.807$ 0.02171 12.24%
Power Breakdown for Hybrid Beamforming with 4 Antennas
Receiver LNA Phase Shifters Combiners Mixers NL Operator ADC[40] Total (mW) Rate Rate/Power % Improvement
Linear Operator $4\times 10$ $8\times 11.5$ $2\times 19.5$ $2\times 14.2$ 00 $2\times 2\times 70fJ\times 2^{8}\times 10^{9}\times 10^{3}$ $271.08$ $16$ $0.05902$
1 Env. Detector $4\times 10$ $8\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 3$ $8\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $214.36$ $11.6147$ 0.05418
3 Env. Detectors $4\times 10$ $8\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 9$ $8\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $226.36$ $15.6147$ 0.06898 16.87%
6 Env. Detectors $4\times 10$ $8\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 18$ $8\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $244.36$ $21.6147$ 0.08845 49.86%
$2^{nd}$ degree Poly. $4\times 10$ $8\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 5$ $8\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $218.36$ $11.6147$ 0.05319
$4^{th}$ degree Poly. $4\times 10$ $8\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 10$ $8\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $228.36$ $13.6147$ 0.05962 1.01%
Power Breakdown for Hybrid Beamforming with 16 Antennas
Receiver LNA Phase Shifters Combiners Mixers NL Operator ADC[40] Total (mW) Rate Rate/Power % Improvement
Linear Operator $16\times 10$ $32\times 11.5$ $2\times 19.5$ $2\times 14.2$ 00 $2\times 2\times 70fJ\times 2^{8}\times 10^{9}\times 10^{3}$ $667.08$ $16$ $0.02399$
1 Env. Detector $16\times 10$ $32\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 3$ $32\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $637.24$ $15.6147$ 0.02450 2.16%
3 Env. Detectors $16\times 10$ $32\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 9$ $32\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $649.24$ $19.6147$ 0.03021 25.96%
6 Env. Detectors $16\times 10$ $32\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 18$ $32\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $667.24$ $25.6147$ 0.03839 60.05%
$2^{nd}$ degree Poly. $16\times 10$ $32\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 5$ $32\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $641.24$ $15.6147$ 0.02435 1.52%
$4^{th}$ degree Poly. $16\times 10$ $32\times 11.5$ $2\times 19.5$ $2\times 14.2$ $2\times 10$ $32\times 2\times 70fJ\times 2^{3}\times 10^{9}\times 10^{3}$ $651.24$ $17.6147$ 0.02705 12.77%
@@
For each RX, a four bit binary weighted current adder is implemented with 1.6mW power consumption
Both power consumption of 3-bit-ADC and eight-bit ADC are normalized from the reference
@@
Figure 8: The breakdown of power consumption for the receiver in a conventional architecture and the proposed architecture.
We further evaluate the achievable rates using the beamforming architecture described in Section 6(c). We numerically simulate the communication system for $n_{r}=2$, $\ell=2$, and $n_{q}\in\{3,4\}$ and find an estimate of the achievable rates empirically ( Figure 6(c)). To compute the rate, we have optimized the parameter $\zeta$ in Equation (15) via brute-force search over the interval [0,1] with step-size 0.01. We have simulated the channel by generating 150000 independent and identically distributed samples of noise vectors and input messages, and have empirically estimated the transition probability of the discrete channel resulting from the quantization process. We have found the optimal probability distribution $P_{X}$ using the Blahut-Arimoto algorithm. It can be observed in Figure 6(c) that for both $n_{q}=3$ and $n_{q}=4$, the high SNR rate is larger than the $n_{r}\left(1+\log{\left(\frac{n_{q}-n_{r}}{n_{r}}\right)}\right)$ lower-bound given in Theorem 5, and it significantly larger than the linear processing baselines.
VIII Circuit Design and Power Analysis